# SOFIA/HAWC+ Far-Infrared Polarimetric Large Area CMZ Exploration Survey. V. The Magnetic Field Strength and Morphology in the Sagittarius C Complex

### v1

@@ 5.2.2 X-rays in the Sgr C H II region
Additional morphological correspondence is observed in the X-ray images of Wang (2021). Figure 15 overlays the 2.5-4 keV Chandra X-ray image from Wang (2021), with the 214 µm pseudovectors.
Figure 15: The 2.5-4 keV Chandra X-ray image reported by Wang (2021). The white contour outlines the [C II] emission in the $-41$ km s-1 velocity channel at the intensity of 13 K (Riquelme et al., In-preparation). The yellow segments are the 214 µm magnetic field pseudovectors made from Paré et al. (2024). The pink-filled star symbols mark the locations of two Wolf-Rayet stars (Geballe et al., 2019). The purple lines denote the location of the Sgr C H II region and the Sgr C NTF, and the green triangle indicates the location of the EGO at the “head” of the head-tail cloud.
Toward the Sgr C H II region and its [C II] shell, much of the X-ray emission appears to arise from within the H II region. The emission becomes fainter in the [C II] shell marked by the white contours in Figure 15. If the X-rays arise from within the HII region, the relative paucity of X-ray emission towards the southeast in the bands between 1 – 6 keV can perhaps be attributed to absorption by the head-tail cloud situated in front of the southeast portion of the H II region.
A somewhat stronger region of extended X-ray emission lies outside and to the southwest of the Sgr C H II region at $l=359.42$, $b=-0.12$. It is possibly connected to the X-ray emission within the [C II] cavity through a “bridge” at $l=359.425$, $b=-0.103$, although it could also be a slightly overlapping, independent source at a different distance. This X-ray feature, clearly external to the Sgr C H II region, exhibits a strong spatial anti-correlation with far-IR emission and with the presence of significant 214 µm magnetic field detections (see Figure 3).
We note again that the location of the X-ray emission to the southwest of the Sgr C HII region coincides in projection with the northeast ends of the G359.33-0.15 radio filaments (displayed in Figure 11 and detailed in Section 5.1.3), though it is not yet clear whether there is any physical association between these two features.
@@
To the southeast of the Sgr C H II region lies the head-tail cloud. The head-tail cloud consists of two parts: an exceptionally bright compact source (the G359.43-0.10 EGO or the “head”; Kendrew et al. 2013; Crowe et al. 2024) and a diffuse and cometary ridge of high far-IR luminosity, high column density (see Figure 6), and low polarization fraction extending towards the southeast (the “tail” addressed in Section 5.1.5). This cloud is also found to exhibit a consistently high turbulent-to-ordered ratio ($\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$), as detailed in Section 3.2.1.Figure 16 shows the magnetic field pseudovectors of the head-tail cloud and the $-65$ km s-1 cloud overlaid with various background images.
Figure 16: The magnetic field pseudovector representation from FIREPLACE coinciding with the head-tail cloud and the $-65$ km s-1 cloud, displayed at wavelengths of 8 µm (Stolovy et al., 2006), 250 µm (Molinari et al., 2010), 850 µm (Pierce-Price et al., 2000), and 20 cm (Heywood et al., 2022). The orange pseudovectors in the lower left panel are the 850 µm magnetic field pseudovectors from Lu et al. (2024). The purple rectangle in the lower right panel denotes the extent of source AGAL 359.474-0.152 detected in the 870 µm ATLASGAL survey (Contreras et al., 2013).
Initially discovered by Forster & Caswell (2000) and Yusef-Zadeh et al. (2009), Kendrew et al. (2013) confirm the existence of early-stage star formation within the EGO, driving multiple jet-like protostellar outflows, which is detailed Crowe et al. (2024). These magnetic field orientations can be due to the cloud being sheared along its long dimension. Alternatively, the head-tail cloud is moving as a unit with constant velocity, and the magnetic field lines threading through that cloud, originally roughly perpendicular to the cloud, can be deformed and dragged into their observed orientations by the bulk motion of the cloud, into which the field lines are frozen. Towards the southeast of the EGO, the tail forms a coherent structure continuous with the $-65$ km s-1 cloud, as depicted in Figure 8.
The JCMT/POL2 polarimetric study surveyed this region at 850 µm (Lu et al. 2024; shown in orange in the lower left panel of Figure 16). We note that the 850 µm magnetic field pseudovector orientations near the “head” generally agree with those of Paré et al. (2024). However, in the body of the $-65$ km s-1 cloud near $l=359.47$, $b=-0.16$, there are several 850 µm magnetic field detections which appear roughly parallel to the Galactic plane and perpendicular to those detected at 214 µm. This misalignment could occur if the polarimetric surveys at different wavelengths are detecting emission from different temperature regimes within the clouds or different components of the same cloud along a given line of sight.
@@
Figure 17: Various morphological features within the $-65$ km s-1 cloud. The yellow segments are the 214 µm magnetic field pseudovectors from Paré et al. (2024), superposed upon the Herschel/SPIRE 250 µm intensity image. The green contours are the $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ ratio at values of (0.1, 0.6, 1.2). The red contours are the 2.5-4 keV X-ray emission from Wang (2021) at (5000, 7000) ${\rm counts\,s^{-1}\,arcmin^{-2}}$. The cyan outlines indicate the extent of the Sgr C H II region and the Sgr C NTF. The purple square denotes the extent of source AGAL 359.474-0.152 detected in the 870 µm ATLASGAL survey (Contreras et al., 2013).
@@
  • Cloud #1. This cloud, centered at $l=359.495$, $b=-0.11$, lies between the Galactic plane to the north (the far-IR-dark region near the top of Figure 16) and the high turbulent-to-ordered ratio “ridge” (the green contours in Figure 16) in the south. Lu et al. (2024) argue that the field morphology in this region, which is parallel to the Galactic plane, is due to its interaction with numerous H II regions along the Galactic plane (Carey et al., 2009; Hankins et al., 2020).
  • Cloud #2. This cloud, centered at $l=359.47$, $b=-0.15$, is located to the south of Cloud #1, bordered by the high turbulent-to-ordered ratio “ridge” to the north and the 2.5-4 keV X-ray emission source to the west (orange contours in Figure 17). The magnetic field within Cloud #2 is uniformly perpendicular to the Galactic plane and to the magnetic field orientation observed in Cloud #1. However, the magnetic field transition from Cloud #1 to Cloud #2 takes place toward the high turbulent-to-ordered ratio “ridge,” thus obscuring the transition from our perspective. We note additionally that the SiO emission in Figure 9 is exclusively found in Cloud #2, likely supporting the case that the magnetic field systems in these listed clouds are segregated.
@@
We note that the observation we provided in this section only pertains to the Paré et al. (2024) 214 µm magnetic field measurements. For instance, we do not take into account the observed JCMT/POL2 pseudovectors in Cloud #2 (Lu et al., 2024). However, we note again that the JCMT/POL2 measurements may be sampling from a different temperature regime or another cloud along the same line of sight. A multi-wavelength dust polarimetry measurement that covers a larger area is warranted to properly interpret the observational differences at the two wavelengths.
@@ 6 Conclusion
This paper presents a detailed analysis of the 214 µm magnetic field pseudovectors from the Paré et al. (2024) dust polarimetry results and their relationship with prominent structures in the Sgr C complex.
@@
  • Through a modified DCF analysis of individual clouds, we found the sky-plane magnetic field strength to be $30~{}\mu{\rm G}-1~{}{\rm mG}$ in Sgr C. Regions with higher turbulence typically have weaker magnetic field strength, such as the head-tail cloud and near the Sgr C H II region. The dominant determinant of the plane-of-sky magnetic field strength is the local turbulence.
  • The expansion of the Sgr C H II region causes extensive turbulent behaviour in its vicinity, as shown by our DCF analysis. We found that the H II region’s interior has been cleared of cold dust content since no far-IR emission is seen there, thereby creating a cavity devoid of significant 214 µm magnetic field pseudovectors. Additionally, we found that the 214 magnetic field orientation is largely tangential to the [C II] shell surrounding the H II region found by Riquelme et al. (In-preparation), and briefly shown in Figure 14. The angular dispersion analysis again confirms the elevated turbulence in this shell, accompanied by existing chemical shock tracers. We note an intriguing diffuse source of X-ray emission to the southwest of the Sgr C H II region, which warrants further study as potentially resulting from past activity in the Sgr C region.
  • The magnetic field morphology near the base of the Sgr C NTF is consistent with the hypotheses that either magnetic field line reconnection or diffusive shock acceleration at the ionization/compression front of the H II region are responsible for the relativistic electrons that illuminate the NTF with their synchrotron emission. The field direction in the $-90$ km s-1 cloud is also found to be parallel to the filament.
  • We have examined the morphological correspondences between the 214 and 850 µm magnetic field pseudovectors and numerous radio features, such as Source C, FIR-4, and the G359.33-0.15 filaments. From these comparisons, we have concluded that many radio structures in the Sgr C region exhibit morphological correspondences with the observed magnetic field direction. Future multi-wavelength dust polarimetry surveys, coupled with detailed molecular line surveys, are needed to elucidate the dynamics of these structures and how the magnetic field and the gas are interacting.
@@ 7 Acknowledgements
This work is primarily based on observations made with the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA). SOFIA was jointly operated by the Universities Space Research Association, Inc. (USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50 OK 2002 to the University of Stuttgart. Financial support for this work was provided by NASA through award #09-0054 through USRA. This work has made extensive use of NASA’s Astrophysics Data System (http://ui.adsabs.harvard.edu/) and the arXiv e-Print service (http://arxiv.org).
@@
The authors at UCLA Department of Physics & Astronomy acknowledge our presence on the traditional, ancestral, and unceded territory of the Gabrielino/Tongva peoples. Much of this work is completed at Meadowridge School, located on the ancestral and unceded territories of the Katzie, Kwantlen, and Coast Salish Peoples, as well as the University of Chicago, located on the ancestral and unceded territories of the Kickapoo, Peoria, Potawatomi, Miami, and Sioux people. We value the opportunity to learn, live, and share research and educational experiences on these traditional lands.
@@ Appendix A.1 Autocorrelation Function of Sgr C Clouds
In this section, we present the details of the autocorrelation function computation and effective cloud depth. We show the shapes of the autocorrelation function calculated for the individual Sgr C clouds in Figure A.1.
Figure A.1: The autocorrelation functions computed from individual Sgr C clouds, as shown in Figure 4. The blue curves are the autocorrelation functions ($\langle\bar{P}^{2}(l)\rangle$), and the shaded region indicate the standard error. The purple horizontal line indicates where $\langle\bar{P}^{2}(l)\rangle=0.5$. The burgundy vertical line indicates the HWHM value of the autocorrelation function, which is equivalent to the effective cloud depth.
We refer the readers to Figure 4 for the extent of individual regions. The cloud depth is positively correlated with the size and uniformity of a region. The autocorrelation functions typically exhibit a uniform decline, except for FIR-4 where a “bump” near $l\approx 1.4$′ is observed (central panel). The shape of the FIR-4 autocorrelation function is likely a result of its size. The vertical extent of FIR-4 is roughly 1.4′ or 0.02, therefore the autocorrelation function likely displays the increase of polarized flux towards the periphery of FIR-4.

### v2

@@ 5.2.2 X-rays in the Sgr C H II region
[…]
Figure 15: The 2.5–4 keV Chandra X-ray image reported by Wang (2021). The white contour outlines the [C II] emission in the $-41$ km s-1 velocity channel at the intensity of 13 K (Riquelme-Vásquez et al., 2025, submitted). The yellow segments are the 214 µm magnetic field pseudovectors from Paré et al. (2024). The pink-filled star symbols mark the locations of two Wolf-Rayet stars (Geballe et al., 2019). The purple shapes denote the location of the Sgr C H II region and the Sgr C NTF, and the green diamond indicates the location of the EGO at the “head” of the head-tail cloud.
This figure shows that the brightest portions of the secondary X-ray peak fit within the [C II] shell surrounding the H II region (white contour, from Riquelme-Vásquez et al. 2025, submitted). Fainter X-ray emission extends away from the H II region both toward the northwest and to the southwest, where it extends continuously at a low level to the brightest X-ray peak lying $\sim 2$’ (5 pc) from the H II region, at $l=359.42$, $b=-0.12$. The two peaks of X-ray emission are possibly connected through a “bridge” at $l=359.425$, $b=-0.103$, although they could also be slightly overlapping, independent sources at different distances.
If the X-rays from the secondary source arise from within the H II region, they could indicate that a supernova has taken place among the presumed cluster of massive stars that powers the H II region. The relative paucity of X-ray emission towards the southeast in the bands between 2.5 – 4 keV can perhaps be attributed to absorption by the head-tail cloud situated in front of the southeast portion of the H II region (Tsuru et al. 2009 also raised this same point to account for the shape of their X-ray contours at this location).
The stronger peak of extended X-ray emission to the southwest of the Sgr C H II region exhibits a strong spatial anti-correlation with far-IR emission and with the presence of significant 214 µm magnetic field detections (see Figure 3).
We note again that the location of the X-ray emission to the southwest of the Sgr C H II region coincides in projection with the northeast ends of the G359.33-0.15 radio filaments (displayed in Figure 11 and detailed in Section 5.1.3), although it is not yet clear whether there is any physical association between these two features.
@@
To the southeast of the Sgr C H II region lies the head-tail cloud. The head-tail cloud consists of two parts: an exceptionally bright compact source (the G359.43-0.10 EGO or the “head”; Kendrew et al. 2013; Crowe et al. 2024; green star in the lower left panel of Figure 16) and a diffuse and cometary ridge of high far-IR luminosity, high column density (see Figure 6), and low polarization fraction extending towards the southeast (the “tail” addressed in Section 5.1.5). This cloud is also found to exhibit a consistently high turbulent-to-ordered ratio ($\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$), as detailed in Section 3.2.1. Figure 16 shows the magnetic field pseudovectors of the head-tail cloud and the $-65$ km s-1 cloud overlaid with various background images.
Figure 16: The magnetic field pseudovector representation from FIREPLACE coinciding with the head-tail cloud and the $-65$ km s-1 cloud, displayed at wavelengths of 8 µm (Stolovy et al., 2006), 250 µm (Molinari et al., 2010), 850 µm (Pierce-Price et al., 2000), and 20 cm (Heywood et al., 2022). The orange pseudovectors in the lower left panel are the 850 µm magnetic field pseudovectors from Lu et al. (2024). The purple rectangle in the lower left panel denotes the extent of source AGAL 359.474-0.152 detected in the 870 µm ATLASGAL survey (Contreras et al., 2013). The white arrow in the lower left panel labels the location of the EGO.
Initially discovered by Forster & Caswell (2000) and Yusef-Zadeh et al. (2009), Kendrew et al. (2013) confirm the existence of early-stage star formation within the EGO, driving multiple jet-like protostellar outflows, recently detailed by Crowe et al. (2024). These magnetic field orientations can be due to the cloud being sheared along its long dimension. Alternatively, the head-tail cloud is moving as a unit with constant velocity, and the magnetic field lines threading through that cloud, originally roughly perpendicular to the cloud, can be deformed and dragged into their observed orientations by the bulk motion of the cloud, into which the field lines are frozen. Towards the southeast of the EGO, the tail forms a coherent structure continuous with the $-65$ km s-1 cloud, as depicted in Figure 8.
The JCMT/POL2 polarimetric study surveyed this region at 850 µm (Lu et al. 2024; shown in orange in the lower left panel of Figure 16). We note that the 850 µm magnetic field pseudovector orientations near the “head” generally agree with those of Paré et al. (2024). However, in the body of the $-65$ km s-1 cloud near $l=359.47$, $b=-0.16$ (purple rectangle in the lower left of Figure 16), there are several 850 µm magnetic field detections which appear roughly parallel to the Galactic plane and perpendicular to those detected at 214 µm. This misalignment could occur if the polarimetric surveys at different wavelengths are detecting emission from different temperature regimes within the clouds or different components of the same cloud along a given line of sight.
@@
Figure 17: Various morphological features within the $-65$ km s-1 cloud. The yellow segments are the 214 µm magnetic field pseudovectors from Paré et al. (2024), superposed upon the Herschel/SPIRE 250 µm intensity image. The green contours are the $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ ratio at values of (0.1, 0.6, 1.2). The red contours depict the 2.5-4 keV X-ray intensity from Wang (2021) at (5000, 7000) ${\rm counts\,s^{-1}\,arcmin^{-2}}$. The cyan outlines indicate the extent of the Sgr C H II region and the Sgr C NTF. The purple rectangle denotes the extent of source AGAL 359.474-0.152 detected in the 870 µm ATLASGAL survey (Contreras et al., 2013). The green star labels the location of the EGO. The orange arrows label the extent of clouds #1, #2, and #4 in Figure 4 used in the DCF analysis in Section 3.2.
@@
  • Cloud #1. This cloud, centred at $l=359.495$, $b=-0.11$, lies between the Galactic plane to the north (the far-IR-dark region near the top of Figure 16) and the high turbulent-to-ordered ratio “ridge” (the green contours in Figure 16) in the south. Lu et al. (2024) argue that the field morphology in this region, which is parallel to the Galactic plane, is due to its interaction with numerous H II regions along the Galactic plane (Carey et al., 2009; Hankins et al., 2020).
  • Cloud #2. This cloud, centred at $l=359.47$, $b=-0.15$, is located to the south of Cloud #1, bordered by the high turbulent-to-ordered ratio “ridge” to the north and the 2.5-4 keV X-ray emission source to the west (orange contours in Figure 17). The magnetic field within Cloud #2 is uniformly perpendicular to the Galactic plane and to the magnetic field orientation observed in Cloud #1. However, the magnetic field transition from Cloud #1 to Cloud #2 takes place toward the high turbulent-to-ordered ratio “ridge,” thus obscuring the transition from our perspective. We note additionally that the SiO emission in Figure 9 is exclusively found in Cloud #2, likely supporting the case that the magnetic field systems in these listed clouds are segregated.
@@
We note that the observation we provided in this section only pertains to the 214 µm magnetic field measurements. For instance, we do not take into account the observed JCMT/POL2 pseudovectors in Cloud #2 (Lu et al., 2024). However, we note again that the JCMT/POL2 measurements may be sampling from a different temperature regime or another cloud along the same line of sight. A multi-wavelength dust polarimetry measurement that covers a larger area is warranted to properly interpret the observational differences at the two wavelengths.
@@ 6 Conclusion
This paper presents a detailed analysis of the FIREPLACE magnetic field measurements from the FIREPLACE dust polarimetry survey (Paré et al., 2024) in the Sgr C complex.
@@
  • Through a modified DCF analysis of individual clouds, we find the sky-plane magnetic field strength to be $30~{}\mu{\rm G}-1~{}{\rm mG}$ in Sgr C. Regions with higher turbulence typically have weaker magnetic field strength, such as the head-tail cloud and near the Sgr C H II region. The dominant determinant of the plane-of-sky magnetic field strength is the local turbulence.
  • The expansion of the Sgr C H II region causes extensive turbulent behaviour in its vicinity, as shown by our DCF analysis. We find that the H II region’s interior has been cleared of cold dust content since no far-IR emission is seen there, thereby creating a cavity devoid of significant 214 µm magnetic field pseudovectors. Additionally, we find that the 214 magnetic field orientation is largely tangential to the [C II] shell surrounding the H II region reported by Riquelme-Vásquez et al. (2025, submitted), and briefly shown in Figure 14. The angular dispersion analysis again confirms the elevated turbulence in this shell, accompanied by existing molecular shock tracers. We note an intriguing diffuse source of X-ray emission to the southwest of the Sgr C H II region, which warrants further study as potentially resulting from past activity in the Sgr C region.
  • The magnetic field morphology near the base of the Sgr C NTF is consistent with the hypotheses that either magnetic field line reconnection or diffusive shock acceleration at the ionization/compression front of the H II region is responsible for the relativistic electrons that illuminate the NTF with their synchrotron emission. Also, the field direction in the $-90$ km s-1 cloud is found to be parallel to the filament.
  • We have examined the morphological correspondences between the 214 and 850 µm magnetic field pseudovectors and numerous radio features, such as Source C, FIR-4, and the G359.33-0.15 filaments. From these comparisons, we conclude that many radio structures in the Sgr C region exhibit morphological correspondences with the observed magnetic field direction. Future multi-wavelength dust polarimetry surveys, coupled with detailed molecular line surveys, are needed to elucidate the dynamics of these structures and to determine how the magnetic field and the gas are interacting.
@@ 7 Acknowledgements
This work is primarily based on observations made with the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA). SOFIA was jointly operated by the Universities Space Research Association, Inc. (USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50 OK 2002 to the University of Stuttgart.
Financial support for this work was provided by NASA through award #09-0054 through USRA. This work has made extensive use of NASA’s Astrophysics Data System (http://ui.adsabs.harvard.edu/) and the arXiv e-Print service (http://arxiv.org). In addition, part of this work is completed using the computing resources provided by the University of Chicago’s Research Computing Center.
@@
The authors at UCLA Department of Physics & Astronomy acknowledge our presence on the traditional, ancestral, and unceded territory of the Gabrielino/Tongva peoples. Much of this work was completed at Meadowridge School, located on the ancestral and unceded territories of the Katzie, Kwantlen, and Coast Salish Peoples, as well as the University of Chicago, located on the ancestral and unceded territories of the Kickapoo, Peoria, Potawatomi, Miami, and Sioux people. We value the opportunity to learn, live, and share research and educational experiences on these traditional lands.
@@ Appendix A.1 Autocorrelation Function of Sgr C Clouds
In this section, we present more details of the autocorrelation function computation and effective cloud depth. We show the shapes of the autocorrelation function calculated for the individual Sgr C clouds in Figure A.1.
Figure A.1: The autocorrelation functions computed from individual Sgr C clouds, as defined in Figure 4. The blue curves are the autocorrelation functions ($\langle\bar{P}^{2}(l)\rangle$), and the shaded region indicate the standard error. The purple horizontal line indicates where $\langle\bar{P}^{2}(l)\rangle=0.5$. The burgundy vertical line indicates the HWHM value of the autocorrelation function, which is equivalent to the effective cloud depth.
We refer the readers to Figure 4 for the location and extent of individual regions. The cloud depth is positively correlated with the size and uniformity of a region. The autocorrelation functions typically exhibit a monotonic decline, except for FIR-4 where a “bump” near $l\approx 1.4$′ is observed (central panel). The shape of the FIR-4 autocorrelation function is likely a result of its morphology. The vertical extent of FIR-4 is roughly 1.4′ or 0.02, therefore the autocorrelation function likely displays the increase of polarized flux towards the northern and southern periphery of FIR-4.